# Protecting Copyrighted Material with Unique Identifiers in Large Language Model Training

### v2

@@ Abstract
A major public concern regarding the training of large language models (LLMs) is whether they abusing copyrighted online text. Previous membership inference methods may be misled by similar examples in vast amounts of training data. Additionally, these methods are often too complex for general users to understand and use, making them centralized, lacking transparency, and trustworthiness. To address these issues, we propose an alternative insert-and-detection methodology, advocating that web users and content platforms employ unique identifiers for reliable and independent membership inference. Users and platforms can create their own identifiers, embed them in copyrighted text, and independently detect them in future LLMs. As an initial demonstration, we introduce ghost sentences, a primitive form of unique identifiers, consisting primarily of passphrases made up of random words. By embedding one ghost sentences in a few copyrighted texts, users can detect its membership using a perplexity test and a user-friendly last-$k$ words test. The perplexity test is based on the fact that LLMs trained on natural language should exhibit high perplexity when encountering unnatural passphrases. As the repetition increases, users can leverage the verbatim memorization ability of LLMs to perform a last-$k$ words test by chatting with LLMs without writing any code. Both tests offer rigorous statistical guarantees for membership inference. For LLaMA-13B, a perplexity test on 30 ghost sentences with an average of 7 repetitions in 148K examples yields a 0.891 ROC AUC. For the last-$k$ words test with OpenLLaMA-3B, 11 out of 16 users, with an average of 24 examples each, successfully identify their data from 1.8M examples.
@@ 1 Introduction
Large language models (LLMs) are pre-trained on vast amounts of data sourced from the Internet, while the service providers of commercial LLMs like ChatGPT [48], Bard [20], and Claude [1] hide the details of the training data from public. This raises concerns that the LLMs may be trained with copyrighted material without permission from the creators [28, 23, 33]. Although some efforts have been made to determine whether a specific example are included in the training data [55, 33, 15], definitive evidence of copyright misuse is still lacking. Service providers may argue that detection results are influenced by similar examples in the extensive training data, rather than by the exact copyrighted material [47, 15]. Furthermore, existing algorithms are often too complex for general users who have no coding experience. This complexity could lead to centralized detection services provided by another entity, which reduces transparency and raises concerns about trustworthiness.
@@
To ensure transparent and trustworthy protection of copyrighted material [^1] Any creative, intellectual, or artistic text presented on the Internet, such as poems, blogs, fiction, and code., we propose an alternative insert-and-detect methodology for general web users and content platforms (e.g., Quora, Medium, Reddit, GitHub). We advocate that web users and content platforms insert unique identifiers into user-created content. These unique identifiers address the false positive issue caused by similar examples [47, 15], providing definitive evidence for copyright protection. The entire process is transparent, allowing users and content platforms to create unique identifiers, embed them in online copyrighted material, and perform detection independently.
To demonstrate the concept, we introduce ghost sentences as a primitive implementation of unique identifiers. A ghost sentence is distinctive because it primarily consists of a randomly generated diceware passphrase [53]. As shown in Figure 1, users or content platforms can insert a ghost sentences, along with customized prefixes, into various online documents. Given an LLM, users can test whether the LLM utilizes their data by performing tests with ther copyrighted examples containing ghost sentences. Based on the repetition of ghost sentences, users can perform perplexity test or user-friendly last-$k$ words test to detect training membership. The basis of the perplexity test is: an LLM trained on natural language should have high perplexity for the passphrases in ghost sentences, as it is a combination of random words. During test, users can generate a new set of ghost sentences, obtain the perplexity distribution, and use the distribution to perform a hypothesis test for membership inference. For a LLaMA-13B model [60], a perplexity test for 30 ghost sentences with a 7 average repetition in 148K examples has a 0.891 ROC AUC.
@@
In potential application scenarios, a user-friendly test method without requiring any expert knowledge is more practical, and we design the last-$k$ words test for general users by simply chatting with LLMs. When the repetition of ghost sentences increases, LLMs are likely to achieve verbatim memorization [9, 10, 27, 28] of the passphrases in ghost sentences. In this case, users can require LLMs to complete the last $k$ words of a ghost sentence, given its preceding context as a prompt like Figure 1 ( a real case at Figure 5). Due to the randomness of passphrases, it is statistically guaranteed that if an LLM can complete the last $k\geq 1$ words, it must have been trained with the ghost sentence. In experiments with an OpenLLaMA-3B [19] model, 11 out 16 users successfully identify their data from the LLM generation. These 16 users have 24 examples with ghost sentences on average and contribute 383 examples to total 1.8M training documents (0.022%).
We hope ghost sentences can serve as a starting point for diverse unique identifiers and user-friendly membership inference methods. A single pattern of unique identifiers is insufficient because it may eventually be filtered out, despite the significant cost of filtering hidden sentences from terabytes or even petabytes of data. Moreover, various identifiers are not mutually exclusive. Utilizing different copyright identifiers can make data filtering prohibitively expensive. Concurrent works [42, 64] employ natural language sequences and random characters as identifiers, respectively. Users can choose one of these identifiers, making filtering difficult. Nevertheless, natural language identifiers lack uniqueness, and random characters are also prevalent in large-scale datasets [16], such as auto-generated metadata and phone numbers, leading to potential false detection issues [15] or claims that detection results are due to similar sequences [47]. Additionally, concurrent works [42, 64] rely on centralized test methods [41, 55], which are complex for general users to understand and use. Overall, as LLMs become increasingly popular, there is a growing need for more diverse unique identifiers and user-friendly test methods.
@@
Figure 1: Insertion and detection of ghost sentences. A custom prefix is added to increase the difficulty of detection. The entire sentence can be made imperceptible on webpages. Given an LLM, users can perform a perplexity test or a last-$k$ words test for reliable membership inference.
@@ 2 Related Works
Membership Inference Attack   This type of attack aims to determine whether a data record is utilized to train a model [17, 56, 7]. Typically, membership inference attacks (MIA) involve observing and manipulating confidence scores or loss of the model [67, 57, 40], as well as training an attack model [56, 25]. Duan et al[15] conduct a large-scale evaluation of MIAs over a suite of LLMs trained on the Pile [18] dataset and find MIAs barely outperform random guessing. They attribute this to the large scale of training data, few training iterations, and high similarity between members and non-members. Shi et al[55] utilize wiki data created after LLMs training to distinguish the members and non-members. Nevertheless, the concern that similar examples in the large-scale training data may lead to ambiguous inference results remains.
Machine-Generated Text Detection   Text watermark [29, 21, 36, 37, 14] aims to embed signals into machine-generated text that are invisible to humans but algorithmically detectable. Generally, LLMs are required not to generate tokens from a red list. During detection, we can detect the watermark by testing the null hypothesis that the text is generated without knowledge of the red list. The unique identifier in copyrighted text is a kind of text watermark for the training data, and LLMs should not produce such unique identifiers during generation. A few other methods [44, 3, 43] try to detect machine-generated text without modifying the generation content. They are mainly based on the assumption that the patterns of log probabilities of human-written and machine-generate text have distinguishable discrepancies.
Training Data Extraction Attack   The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim [8, 27, 69]. Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs [10, 46, 31, 30]. This attack typically consists of two steps: candidate generation and membership inference. The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM. Carlini et al[8] quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples. Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more. Karamolegkou et al[28] demonstrate that LLMs can achieve verbatim memorization of literary works and educational material. We also provide an similar example in Figure 5 in Appendix C.
@@ 3.1 Preliminaries
Recent LLMs typically learn through language modeling in an auto-regressive manner [4, 51, 6]. For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$. During training, LLMs are optimized to maximize the joint probability of $x$: $p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$
@@
Creation of Ghost Sentences   The main part of a ghost sentence is a diceware passphrase [53]. Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$. $V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times. For a diceware passphrase with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of indexed webpages estimated by worldwidewebsize.com. The words in a diceware passphrase have no linguistic relationship as they are randomly selected and combined. Users can customize ghost sentences by add prefixes to passphrases as shown in Figure 1. It is recommended to use passphrases with more than 8 words and insert ghost sentences in the latter half of a document. We provide a few examples of ghost sentences in Appendix I.
Statistics of Users on Reddit   In Appendix D, we provide the statistics of users in Webis-TLDR-17 [62], a subset of Reddit data contains 3.8M examples from 1.4M users. The distribution of the number of document per user is long-tailed. Users with more than 4 and 9 examples contribute 41% and 22% of all data, respectively. These users can insert ghost sentences by themselves, other users contribute about 60% examples may need assistance from the content platform.
Null Hypothesis   We detect ghost sentences by testing the following null hypothesis,
@@
3.2 Perplexity Test
The perplexity of a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ given context ${c}=(s_{1},\ldots,s_{j})$ is:
For simplicity, we only consider the perplexity of passphrases in ghost sentences. Passphrases are combinations of random words. If the null hypothesis is true, the LLM is basically doing random guessing given a vocabulary with size $V$, and the value of $\texttt{PPL}(g)$ should be very high.
(a) OpenLLaMA-3B
(b) LLaMA-7B
(c) LLaMA-13B
Figure 2 presents the perplexity discrepancy between normal context ($\texttt{PPL}(c)$) and ghost sentences ($\texttt{PPL}(g)$ given $c$). On average, the perplexity of ghost sentences are much higher than that of natural language. Given an LLM, a ghost sentence $g$, and a context $c$, we can use the empirical perplexity distribution of ghost sentences (unseen by the LLM) in Figure 2 to perform a hypothesis test. If $\texttt{PPL}(g)$ is smaller than the critical value at a certain significant level, we will reject the null hypothesis $H_{0}$. For example, if $\texttt{PPL}(g)<157$ for a LLaMA-7B model in Figure 2, we will reject $H_{0}$ and the probability of a false positive is less than 1%. Perplexity test requires one ghost sentences to repeat for a few times in the training data of LLMs. For a LLaMA-13B model fine-tuned on 148K examples with 30 ghost sentences repeat 5 times on average, a perplexity test can achieve 0.393 recall with a significant level 0.05 after 1 epoch fine-tuning. The recall increase to 0.671 if the average repetition becomes 7.
@@ 3.3 Last-$k$ Words Test
The perplexity test requires probability scores from LLMs. While some LLM service providers, such as OpenAI, provide APIs for accessing probability scores, it can be challenging for general users to perform perplexity test on their own. In the following content, we demonstrate that users can perform a hypothesis test using solely the generation function of LLMs. However, this approach requires the ghost sentence to be repeated more times compared to a perplexity test.
@@
Suppose the length of passphrase $q=10$ and $V^{\star}=7,776$, with $n_{g}=1$. This results in a z-score of 27.85. The critical value at a significance level of 0.01 is only 2.58. In this case, we reject the null hypothesis, and the probability of a false positive is nearly 00. In practice, as ghost sentences in the training data increase, $1/V^{\star}$ also increases, and a large $n_{g}$ may be required for the test. When $n_{g}=2$, the test can reject the null hypothesis even if $251/V^{\star}=1/$ 25 at a significant level $0.01$. A probability $1/25$ is clearly not normal for generating random words. Our analysis for ghost sentence detection is similar to that for detecting text watermark [29].
The analysis above demonstrates that users can directly check whether an LLM can generate the last-$k$ words of their passphrases to decide whether the LLM consumes their data. $k=1$ or $k=2$ can already guarantee the robustness of ghost sentence detection. To understand how many repetition times for ghost sentences are required for the last-$k$ words test, we define two quantitative metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc):

### v3

@@ Abstract
A primary concern regarding training large language models (LLMs) is whether they abuse copyrighted online text. With the increasing training data scale and the prevalence of LLMs in daily lives, two problems arise: 1) false positive membership inference results misled by similar examples; 2) membership inference methods are usually too complex for end users to understand and use. To address these issues, we propose an alternative insert-and-detect methodology, advocating that web users and content platforms employ unique identifiers for reliable and independent membership inference. Users and platforms can create their identifiers, embed them in copyrighted text, and independently detect them in future LLMs. As an initial demonstration, we introduce ghost sentences and a user-friendly last-$k$ words test, allowing end users to chat with LLMs for membership inference. Ghost sentences consist primarily of unique passphrases of random natural words, which can come with customized elements to bypass possible filter rules. The last-$k$ words test requires a significant repetition time of ghost sentences ($\geq 10$). For cases with fewer repetitions, we designed an extra perplexity test, as LLMs exhibit high perplexity when encountering unnatural passphrases. We also conduct a comprehensive study on the memorization and membership inference of ghost sentences, examining factors such as training data scales, model sizes, repetition times, insertion positions, wordlist of passphrases, alignment, etc. Our study shows the possibility of applying ghost sentences in real scenarios and provides instructions for the potential application.
@@ 1 Introduction
Large language models (LLMs) are pre-trained on vast amounts of data sourced from the Internet, while the providers of commercial LLMs like ChatGPT, Bard, and Claude do not disclose the details of the training data. This raises concerns that LLMs may be trained with copyrighted material without permission from the creators (Karamolegkou et al., 2023; Henderson et al., 2023; Li et al., 2024). Some efforts have been made to determine whether a specific example is included in the training data (Mattern et al., 2023; Meeus et al., 2024; Shi et al., 2024; Li et al., 2024). However, the false positive membership inference results caused by similar examples are common (Duan et al., 2024). Service providers might argue that detection results could be confused by similar examples in massive data rather than the exact copyrighted content (OpenAI, 2019). Additionally, these membership inference methods are often too complex for end users without coding experience or expert knowledge. This complexity could lead to centralized detection services, which reduce transparency and raise concerns about trustworthiness.
@@
For transparent and reliable protection of copyrighted material [^1] Any creative, intellectual, or artistic text presented on the Internet, such as poems, blogs, fiction, and code., we propose an alternative insert-and-detect methodology for general web users and content platforms (e.g., Quora, Medium, Reddit, GitHub). We advocate that web users and content platforms insert unique identifiers into copyrighted content. These identifiers help address the issue of false positives caused by similar examples (OpenAI, 2019; Duan et al., 2024), providing definitive evidence for copyright protection. The process should be transparent, allowing users and content platforms to create unique identifiers, embed them in online copyrighted material, and perform detection independently.
@@
To demonstrate the concept, we introduce ghost sentences as a primitive implementation of unique identifiers, as well as a user-friendly last-$k$ words test for their membership inference. A ghost sentence is distinctive because it primarily consists of a randomly generated diceware passphrase (Reinhold, 1995). As shown in Figure 1, users or content platforms can insert a ghost sentence, along with customized elements, into various online documents. When the repetition of ghost sentences increases, LLMs are likely to achieve verbatim memorization (Carlini et al., 2019, 2021; Ishihara, 2023; Karamolegkou et al., 2023) of the passphrases in ghost sentences. In this case, users can prompt LLMs to complete the last $k$ words of a ghost sentence, using the preceding context, as shown in Figure 1. For example, the last-$k$ word test can be performed on ChatGPT using content from popular books, as demonstrated in Figure 5. Due to the randomness of passphrases, it is statistically guaranteed that if an LLM can complete the last $k\geq 1$ words, it must have been trained with the ghost sentence. In experiments with an OpenLLaMA-3B (Geng and Liu, 2023) model, 11 out 16 users successfully identify their data from the LLM generation. These 16 users have 24 examples with ghost sentences on average and contribute 383 examples to a total of 1.8M training documents. Ghost sentences account for only 0.0017% of all training tokens.
The last $k$ words test is user-friendly but requires a non-trivial repetition time of ghost sentences ($\geq 10$). Following previous membership inference methods based on loss, entropy, or probability of predictions (Yeom et al., 2018; Carlini et al., 2021; Shi et al., 2024), we design an alternative perplexity test for less frequently repeated ghost sentences. An LLM trained with natural languages should exhibit high perplexity for the passphrase in a ghost sentence, as it consists of random words. During the perplexity test, users can generate a new set of ghost sentences, obtain the perplexity distribution, and use the distribution to perform a hypothesis test for membership inference. For a LLaMA-13B model (Touvron et al., 2023a), a perplexity test for 30 ghost sentences, with an average repetition of 7 in 148K examples, achieves a 0.891 ROC AUC.
@@
Figure 1: Insertion and test of ghost sentences. A ghost sentence primarily consists of a unique passphrase, with customizable elements like a prefix added to bypass potential filters. Given an LLM, users can conduct a last-$k$ words test by interacting with the LLM for reliable membership inference. Alternatively, users can perform a perplexity test if prediction scores are available.
We also comprehensively study different factors influencing the memorization and membership inference results of ghost sentences. A few key observations are as follows: 1) The memorization of ghost sentences is jointly decided by their quantity and average repetition. Ghost sentences with a word length $\geq 8$, an average repetition $\geq 5$, and a proportion $\geq 0.0016\%$ of training tokens are recommended. 2) It is better to insert ghost sentences in the latter half of a document. 3) A curated wordlist for the generation of passphrases is necessary. We suggest using a well-maintained wordlist from the Electronic Frontier Foundation. 4) Further model alignment (Ouyang et al., 2022; Rafailov et al., 2023) will not affect the memorization of ghost sentences. 5) The larger the model, the smaller the repetition times for memorization. This is consistent with previous works (Carlini et al., 2023). Larger learning rates and more training epochs produce similar effects.
A single pattern of unique identifiers is insufficient, as it may eventually be filtered out, despite the significant cost of filtering hidden sentences from terabytes or even petabytes of data. As LLMs become increasingly popular in daily lives, there is a growing need for diverse unique identifiers and user-friendly test methods. Different copyright identifiers are not mutually exclusive and can work together to make filtering intractable. Wei et al. (2024) adopt random characters, which also qualify as unique identifiers. Nevertheless, relying solely on long random characters risks filter through measures like regular expression matching and semantic checking. Additionally, random characters, such as auto-generated metadata, are prevalent in large-scale datasets (Elazar et al., 2024), which can lead to false detection issues (Duan et al., 2024). They also lack a user-friendly membership inference method for end users. We hope ghost sentences can serve as a starting point for creating diverse unique identifiers and user-friendly membership inference methods.
@@ 2 Related Works
Membership Inference Attack
This type of attack aims to determine whether a data record is utilized to train a model (Fredrikson et al., 2015; Shokri et al., 2017; Carlini et al., 2022). Typically, membership inference attacks (MIA) involve observing and manipulating confidence scores or loss of the model (Yeom et al., 2018; Song and Shmatikov, 2019; Mattern et al., 2023), as well as training an attack model (Shokri et al., 2017; Hisamoto et al., 2020). Duan et al(Duan et al., 2024) conduct a large-scale evaluation of MIAs over a suite of LLMs trained on the Pile (Gao et al., 2020) dataset and find MIAs barely outperform random guessing. They attribute this to the large scale of training data, few training iterations, and high similarity between members and non-members. Shi et al(Shi et al., 2024) utilize wiki data created after LLMs training to distinguish the members and non-members. Nevertheless, the concern that similar examples in the large-scale training data may lead to ambiguous inference results remains.
Machine-Generated Text Detection
Text watermark (Kirchenbauer et al., 2023; Gu et al., 2024; Liu et al., 2024, 2023; Ding et al., 2024) aims to embed signals into machine-generated text that are invisible to humans but algorithmically detectable. Generally, LLMs are required not to generate tokens from a red list. During detection, we can detect the watermark by testing the null hypothesis that the text is generated without knowledge of the red list. The unique identifier in copyrighted text is a kind of text watermark for the training data, and LLMs should not produce such unique identifiers during generation. A few other methods (Mitchell et al., 2023; Bao et al., 2024; Mireshghallah et al., 2024) try to detect machine-generated text without modifying the generation content. They are mainly based on the assumption that the patterns of log probabilities of human-written and machine-generate text have distinguishable discrepancies.
Training Data Extraction Attack
The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim (Carlini et al., 2023; Ishihara, 2023; Zhang et al., 2023). Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs (Carlini et al., 2021; Nasr et al., 2023; Lee et al., 2023; Kudugunta et al., 2023). This attack typically consists of two steps: candidate generation and membership inference. The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM. Carlini et al(Carlini et al., 2023) quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples. Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more. Karamolegkou et al(Karamolegkou et al., 2023) demonstrate that LLMs can achieve verbatim memorization of literary works and educational material. We also provide an similar example in Figure 5 in §App. C.
@@ 3.1 Preliminaries
Recent LLMs typically learn through language modeling in an auto-regressive manner (Bengio et al., 2003; Radford et al., 2019; Brown et al., 2020). For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$. During training, LLMs are optimized to maximize the joint probability of $x$: $p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$
@@
Creation of Ghost Sentences
The main part of a ghost sentence is a diceware passphrase (Reinhold, 1995). Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$. $V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times. For a diceware passphrase with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of indexed webpages estimated by worldwidewebsize.com (at least 2.37 billion indexed pages, October, 2024). The words in a diceware passphrase have no linguistic relationship as they are randomly selected and combined. Users can customize ghost sentences by add prefixes to passphrases as shown in Figure 1. It is recommended to use passphrases with more than 8 words and insert ghost sentences in the latter half of a document. We provide a few examples of ghost sentences in §App. G.
Statistics of Users on Reddit
In §App. D, we provide the statistics of users in Webis-TLDR-17 (Völske et al., 2017), a subset of Reddit data contains 3.8M examples from 1.4M users. The distribution of the number of document per user is long-tailed. Users with more than 4 and 9 examples contribute 41% and 22% of all data, respectively. These users can insert ghost sentences by themselves, other users contribute about 60% examples may need assistance from the content platform.
Null Hypothesis
We detect ghost sentences by testing the following null hypothesis,
@@ 3.2 Last-$k$ Words Test
Suppose the length of passphrase $q=10$ and $V^{\star}=7,776$, with $n_{g}=1$. This results in a z-score of $27.85\gg 2.58$; the latter is at a significant level of 0.01. In this case, we reject the null hypothesis, and the probability of a false positive is nearly 00. In practice, as ghost sentences in the training data increase, $1/V^{\star}$ also increases, and a large $n_{g}$ may be required for the test. When $n_{g}=2$, the test can reject the null hypothesis even if $251/V^{\star}=1/$ 25 at a significant level $0.01$. A probability $1/25$ is clearly not normal for generating random words. Our analysis for ghost sentence detection is similar to that for detecting text watermark (Kirchenbauer et al., 2023).
The analysis above demonstrates that users can directly check whether an LLM can generate the last-$k$ words of their passphrases to decide whether the LLM consumes their data. $k=1$ or $k=2$ can already guarantee the robustness of test results. To understand how many repetition times for ghost sentences are required for the last-$k$ words test, we define two quantitative metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc):
@@
3.3 Perplexity Test
The last $k$ words test is user-friendly but requires a significant repetition time  ($>10$) to achieve verbatim memorization of ghost sentences. Inspired by previous membership inference methods based on loss, entropy, or probability of predictions (Yeom et al., 2018; Carlini et al., 2021; Shi et al., 2024), we design a perplexity test for less repeated ghost sentences. The perplexity of a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ given context ${c}=(s_{1},\ldots,s_{j})$ is:
For simplicity, we only consider the perplexity of passphrases, excluding customized elements. Passphrases are combinations of random words. If the null hypothesis is true, the LLM is basically doing random guessing given a vocabulary $V$, and the value of $\texttt{PPL}(g)$ should be high.
(a) OpenLLaMA-3B
(b) LLaMA-7B
(c) LLaMA-13B
Figure 2 presents the perplexity discrepancy between normal context ($\texttt{PPL}(c)$) and ghost sentences ($\texttt{PPL}(g)$ given $c$). On average, the perplexity of ghost sentences are much higher than that of natural language. Given an LLM, a ghost sentence $g$, and a context $c$, we can use the empirical perplexity distribution of ghost sentences (unseen by the LLM) to perform a hypothesis test. If $\texttt{PPL}(g)$ is smaller than the critical value at a certain significant level, we will reject the null hypothesis $H_{0}$. For example, if $\texttt{PPL}(g)<157$ for a LLaMA-7B model in Figure 2, we will reject $H_{0}$ and the probability of a false positive is less than 1%. The perplexity test requires one ghost sentence to be repeated a few times in the training data of LLMs. For a LLaMA-13B model fine-tuned on 148K examples with 30 ghost sentences repeat 5 times on average, a perplexity test can achieve 0.393 recall with a significant level 0.05 after 1 epoch fine-tuning. The recall increases to 0.671 if the average repetition becomes 7.